# Renovating Names in Open-Vocabulary Segmentation Benchmarks

### v1

@@ 0.B.2 More details on the human verification process
Figure 0.B.1: Human verification for upgrading benchmarks. We ask 5 human annotators to verify the names of segments in the validation sets of ADE20K and Cityscapes. As shown in (a), humans are given the top 3 suggestions from the model and are asked to verify the selected name or choose a more matching name from either the other top 3 names or the rest of the candidate names. (b) shows that RENOVATE name suggestions significantly speed up the human verification/annotation process.
@@
Figure 0.B.2: Typical cases when humans and models disagree during human verification. We show two typical cases when humans may choose different names from our model-selected names. (a) shows when segments are too small to recognize, it can be difficult to decide which name is correct. This may indicate a limitation of our renaming method on extremely small objects. (b) shows when there is a lack of sufficient visual cues to infer the names of the segments, the choices are ambiguous. For example, both “river” and “lake” are reasonable choices for the segment in the top right image without more information of the scene. Note (a) and (b) consist only a minority of all the segments.
0.B.3 More details on the open metrics in upgraded benchmarks
@@
0.B.4 Other implementation details
More details on the human study. The human study (Sec. 4.1) is conducted on ADE20K and Cityscapes datasets, which consists of 30249 and 14670 segments in the validation set, respectively. We randomly select 10% of the validation set for human study. The average time for human experts to finish the preference study is about 2 hours.
More details on the renaming model. For the backbone of the renaming model, we use ConvNeXt-Large CLIP [32, 13] from OpenCLIP [13] pretrained on LAION-2B dataset [38]. Following Mask2Former [4], our transformer decoder consists of nine blocks and we use AdamW [23] optimizer with a learning rate of $10^{-4}$ and a multi-step decay schedule with weight decay of 0.05.
@@ 0.B.4 Other implementation details
Appendix 0.C More experiment results
ADE20K [57] Cityscapes [7]
Benchmark Test Name Metric #classes PQ AP mIoU #classes PQ AP mIoU
Original Original Standard 150 21.88 13.94 29.16 19 39.72 27.73 49.60
Open [58] 150 26.31 14.49 34.39 19 44.09 27.91 58.20
OpenSeg Standard 150 21.63 13.95 28.88 19 43.26 28.45 54.53
Open [58] 150 28.02 15.32 38.86 19 46.66 28.88 58.61
Upgraded RENOVATE Standard 578 3.15 1.21 5.64 108 5.93 4.54 6.25
Open [58] 578 10.35 4.03 22.13 108 18.27 15.15 20.20
Table 0.C.1: Open-vocabulary segmentation on upgraded benchmarks using ODISE [48]. Similar to Tab. 2, the upgraded benchmarks are considerably more challenging and the gap between standard and open metrics is significantly larger on the upgraded benchmark, highlighting the need of open metrics for more human-aligned evaluation when names are more similar.
ADE20K [57] Cityscapes [7]
Benchmark Test Name Metric #classes PQ AP mIoU #classes PQ AP mIoU
Original Original Standard 150 23.46 12.80 30.32 19 33.78 18.11 45.35
Open [58] 150 26.53 13.01 36.33 19 37.94 18.35 55.75
OpenSeg Standard 150 23.70 12.84 31.17 19 35.05 17.79 46.46
Open [58] 150 28.04 13.89 39.40 19 38.83 18.62 52.11
Upgraded RENOVATE Standard 578 6.36 3.78 3.90 108 6.12 4.58 5.53
Open [58] 578 24.27 10.78 18.13 108 16.50 15.42 18.89
Table 0.C.2: Open-vocabulary segmentation on upgraded benchmarks using MasQCLIP [51].
More results on upgraded benchmarks. In addition to FC-CLIP model [52] (Sec. 4.2), we further evaluate two other models on the upgraded benchmarks of ADE20K and Cityscapes and compare with the evaluation results on the original benchmarks. The results in Tab. 0.C.1 and Tab. 0.C.2 are consistent with our findings in Sec. 4.2. Specifically, since names in our upgraded benchmarks are more fine-grained and more similar, our benchmarks are more challenging and the discrepancy between standard and open metrics is much greater.
More results on training with renovated names. We further evaluate the models trained with different sets of names using open metrics in Tab. 0.C.3. The results are consistent with Tab. 3 and demonstrate that training with RENOVATE names improves the model’s open-vocabulary generalization ability.
Source dataset Target datasets
MS COCO ADE20K Cityscapes
Train names PQ AP mIoU PQ AP mIoU PQ AP mIoU
Original names 53.55 42.69 64.81 28.24 16.70 38.34 47.57 26.94 63.13
OpenSeg names 55.05 43.47 64.52 30.55 17.84 42.18 46.93 27.51 62.48
RENOVATE 57.34 46.44 66.47 30.89 18.10 44.24 47.90 30.14 65.35
Table 0.C.3: Evaluating models trained with different names using open metrics. Except for the metric, the evaluation process is the same as in Sec. 4.3.
Original name Context names Candidate names Renovated names mIoU gain earth, ground building, tree, stand, water, person, floor, sky, road, snow, car ground cover,earth floor,ground cover, ground,earth surface,snow ground, land,natural ground,outdoor ground,terrain, earth,dirt ground,terrestrial terrain ground cover, terrestrial terrain, natural ground, snow ground, dirt ground, ground, grass +11.50 field lush, field, sky, green, grass, tree, road, hillside, grassy, rural rural field, roadside field, green field, crop field, sports field, grassland, grassy hillside, rural field, roadside field, grassland, crop field, sports field +15.02 rock, stone stone, water, tree, lush, sea, sky, shoreline, coast, creek, stand rock,stone,standing stone,beach pebble, shoreline rock,national park boulders, river stone,stones,lush stone,rocks river stone, stones, beach pebble rock, national park boulders, shoreline rock +8.95 fountain city, fountain, building, garden, square, park, water, hotel, place, design garden fountains,fountain,urban fountains, square fountains,park fountains, water feature designs,hotel water features hotel water features, urban fountains,park fountains +31.58 counter room, area, lobby, reception, shop, store, hotel, design, office, interior checkout counter,reception desk, office counter area,countertop,hotel lobby counter, cashier counter,counter,retail store counter, service desk,storefront counter,information kiosk information kiosk, reception desk, cashier counter, counter, office counter area, checkout counter +20.72 stairway, staircase building, stair, floor, room, ceiling, lead to, pillar, balustrade, stairwell, rail balustrade staircase,ceiling stairwell,stairway, building stairs,staircase,pillar-supported stairs, room stairway,railing stairway,floor staircase pillar-supported stairs,railing stairway, room stairway, floor staircase +14.96 bar bar, restaurant, room, area, hotel, interior, pub, table, house, counter bar table,bar area,restaurant bar,hotel bar, home bar,bar counter,bar interior,bar,pub bar bar area, bar table bar counter +18.58 canopy room, bed, bedroom, house, idea, design, palace, garden, home, royal home canopy,canopy,house canopy, garden canopy,royal palace canopy,patio canopy, bedroom canopy,room canopy bedroom canopy, patio canopy, garden canopy +9.88 grandstand, covered stand stadium, game, field, arena, world, park, football, school, sport, high high-capacity grandstand, park bleachers, covered game stand, world-class viewing area, spectator stand, football field seating,grandstand, sports arena stand,stadium grandstand, school spectator seats sports venue grandstand, sports arena stands, park bleachers, spectator stand football field seating +29.50 swimming pool, swimming bath, natatorium pool, swimming, house, hotel, villa, resort, beach, indoor, luxury, apartment luxury resort pool,swimming bath,villa pool, apartment complex pool,hotel swimming pool, home swimming pool,swimming pool, beachfront swimming pool,indoor swimming bath hotel swimming pool, beachfront swimming pool, villa pool, indoor swimming pool +30.12
Table 0.C.4: Examples of context names, candidate names, and RENOVATE names for classes from ADE20K.
Figure 0.C.1: More examples on name selection based on IoU scores.
Appendix 0.D Examples of context, candidate, renovated names, and name selection
@@
In Tab. 0.C.4, we show examples of original names, context names, candidate names, RENOVATE names, and their mIoU gain in the open-vocabulary segmentation evaluation as in Sec. 4.4. For RENOVATE, the names for each class are simply the set of unique names for all the instances in that class after renaming. We can see that the context names are important for understanding the vague class names such as “earth” and “field” and are helpful for providing more contexts to names with multiple potential meanings like “counter” and “bar”. We can also see that RENOVATE names provide more comprehensive descriptions of the corresponding classes. Using these names to prompt open-vocabulary models leads to significant performance gains, suggesting that our renovated names match better with the visual segments.
@@
In Fig. 0.C.1, we further show more visual examples of the name selection process (Sec. 3.3). Together with Fig. 3 in the main paper, these examples demonstrate that our renaming model provides a meaningful ranking of the names to match with each segment.
Appendix 0.E More qualitative analysis
0.E.1 RENOVATE can refine original names
@@
In Fig. 0.E.1 and Fig. 0.E.2, we present more examples of RENOVATE names along with the original names of segments from ADE20K and MS COCO. We show that RENOVATE refines the original names by being more precise and closer to human natural language.
@@
Figure 0.E.1: More examples of renovated names on segments from the validation set of ADE20K. For each segment, we show the original name below the image and the renovated name in the text box.
@@
Figure 0.E.2: More examples of renovated names on segments from the validation set of MS COCO. For each segment, we show the original name below the image and the renovated name in the text box.
0.E.2 Other potential use cases of RENOVATE
@@
RENOVATE can correct annotation errors. From Fig. 0.E.3, we can see that our renaming model discovers wrong annotations and finds more accurate names through RENOVATE. It’s important to highlight that these renovated names do not belong to the original ground-truth class; instead, our model ranks them over names from the original class. This demonstrates the renaming model’s utility in detecting and correcting inaccuracies in annotations by effectively matching names with visual content.
Figure 0.E.3: RENOVATE can find wrong annotations and suggest corrections.
@@
RENOVATE uncovers segments with shared semantic concepts across datasets. Since RENOVATE aims to describe visual segments using more human-aligned language, it effectively standardizes the naming conventions for identical concepts in different datasets. In Fig. 0.E.4, we demonstrate this with examples from MS COCO and ADE20K. While renaming each dataset separately, we observe that RENOVATE inadvertently reveals the shared name spaces by identifying more appropriate names in natural language to describe the visual contents. This property can be potentially used for analyzing the differences between datasets, merging datasets to construct new benchmarks, and various other data curation processes.
@@
Figure 0.E.4: RENOVATE uncovers segments with shared semantic concepts across datasets.

### v2

@@ C.2 More details on the human verification process
Figure C.1: Human verification for upgrading benchmarks. We ask 5 human annotators to verify the names of segments in the validation sets of ADE20K and Cityscapes. As shown in (a), humans are given the top 3 suggestions from the model and are asked to verify the selected name or choose a more matching name from either the other top 3 names or the rest of the candidate names. (b) shows that RENOVATE name suggestions significantly speed up the human verification/annotation process.
@@
Figure C.2: Typical cases when humans and models disagree during human verification. We show two typical cases when humans may choose different names from our model-selected names. (a) shows when segments are too small to recognize, it can be difficult to decide which name is correct. This may indicate a limitation of our renaming method on extremely small objects. (b) shows when there is a lack of sufficient visual cues to infer the names of the segments, the choices are ambiguous. For example, both “river” and “lake” are reasonable choices for the segment in the top right image without more information of the scene. Note (a) and (b) consist only a minority of all the segments.
C.3 More details on the implementation of open metrics
@@
C.4 Other implementation details
More details on the renaming model. For the backbone of the renaming model, we use ConvNeXt-Large CLIP [23, 62] from OpenCLIP [62] pretrained on LAION-2B dataset [63]. Following Mask2Former [31], our transformer decoder consists of nine blocks and we use AdamW [64] optimizer with a learning rate of $10^{-4}$ and a multi-step decay schedule with weight decay of 0.05. Training one renaming model on 4 A-100 GPUs requires approximately 3 days.
More details on training with renovated names.. We follow the training schedule in prior works [25] to train FCCLIP and use negative sampling with $C=100$ in training all models. Training one model on 4 A-100 GPUs requires approximately 4 days.
@@ C.4 Other implementation details
Appendix D More Name Examples
@@
In Table D.1, we show examples of original names, context names, candidate names, and RENOVATE names. For RENOVATE, the names for each class are simply the set of unique names for all the instances in that class after renaming. We can see that the context names are important for understanding the vague class names such as “earth” and “field” and are helpful for providing more contexts to names with multiple potential meanings like “counter” and “bar”. We can also see that RENOVATE names provide more comprehensive descriptions of the corresponding classes.
Original name Context names Candidate names Renovated names earth, ground building, tree, stand, water, person, floor, sky, road, snow, car ground cover,earth floor,ground cover, ground,earth surface,snow ground, land,natural ground,outdoor ground,terrain, earth,dirt ground,terrestrial terrain ground cover, terrestrial terrain, natural ground, snow ground, dirt ground, ground, grass field lush, field, sky, green, grass, tree, road, hillside, grassy, rural rural field, roadside field, green field, crop field, sports field, grassland, grassy hillside, rural field, roadside field, grassland, crop field, sports field rock, stone stone, water, tree, lush, sea, sky, shoreline, coast, creek, stand rock,stone,standing stone,beach pebble, shoreline rock,national park boulders, river stone,stones,lush stone,rocks river stone, stones, beach pebble rock, national park boulders, shoreline rock fountain city, fountain, building, garden, square, park, water, hotel, place, design garden fountains,fountain,urban fountains, square fountains,park fountains, water feature designs,hotel water features hotel water features, urban fountains,park fountains counter room, area, lobby, reception, shop, store, hotel, design, office, interior checkout counter,reception desk, office counter area,countertop,hotel lobby counter, cashier counter,counter,retail store counter, service desk,storefront counter,information kiosk information kiosk, reception desk, cashier counter, counter, office counter area, checkout counter stairway, staircase building, stair, floor, room, ceiling, lead to, pillar, balustrade, stairwell, rail balustrade staircase,ceiling stairwell,stairway, building stairs,staircase,pillar-supported stairs, room stairway,railing stairway,floor staircase pillar-supported stairs,railing stairway, room stairway, floor staircase bar bar, restaurant, room, area, hotel, interior, pub, table, house, counter bar table,bar area,restaurant bar,hotel bar, home bar,bar counter,bar interior,bar,pub bar bar area, bar table bar counter canopy room, bed, bedroom, house, idea, design, palace, garden, home, royal home canopy,canopy,house canopy, garden canopy,royal palace canopy,patio canopy, bedroom canopy,room canopy bedroom canopy, patio canopy, garden canopy grandstand, covered stand stadium, game, field, arena, world, park, football, school, sport, high high-capacity grandstand, park bleachers, covered game stand, world-class viewing area, spectator stand, football field seating,grandstand, sports arena stand,stadium grandstand, school spectator seats sports venue grandstand, sports arena stands, park bleachers, spectator stand football field seating swimming pool, swimming bath, natatorium pool, swimming, house, hotel, villa, resort, beach, indoor, luxury, apartment luxury resort pool,swimming bath,villa pool, apartment complex pool,hotel swimming pool, home swimming pool,swimming pool, beachfront swimming pool,indoor swimming bath hotel swimming pool, beachfront swimming pool, villa pool, indoor swimming pool
Table D.1: Examples of context names, candidate names, and RENOVATE names for classes from ADE20K.
Figure D.1: More examples on name selection based on IoU scores.
@@
In Fig. D.1, we further show more visual examples of the name selection process (Section 3.3). Together with Fig. 3 in the main paper, these examples demonstrate that our renaming model provides a meaningful ranking of the names to match with each segment.
Appendix E More Qualitative Analysis
E.1 RENOVATE can refine original names
@@
In Fig. E.1 and Fig. E.2, we present more examples of RENOVATE names along with the original names of segments from ADE20K and MS COCO. We show that RENOVATE refines the original names by being more precise and closer to human natural language.
@@
Figure E.1: More examples of renovated names on segments from the validation set of ADE20K. For each segment, we show the original name below the image and the renovated name in the text box.
@@
Figure E.2: More examples of renovated names on segments from the validation set of MS COCO. For each segment, we show the original name below the image and the renovated name in the text box.
E.2 Other potential use cases of RENOVATE
@@
RENOVATE can correct annotation errors. From Fig. E.3, we can see that our renaming model discovers wrong annotations and finds more accurate names through RENOVATE. It’s important to highlight that these renovated names do not belong to the original ground-truth class; instead, our model ranks them over names from the original class. This demonstrates the renaming model’s utility in detecting and correcting inaccuracies in annotations by effectively matching names with visual content.
Figure E.3: RENOVATE can find wrong annotations and suggest corrections.
@@
RENOVATE uncovers segments with shared semantic concepts across datasets. Since RENOVATE aims to describe visual segments using more human-aligned language, it effectively standardizes the naming conventions for identical concepts in different datasets. In Fig. E.4, we demonstrate this with examples from MS COCO and ADE20K. While renaming each dataset separately, we observe that RENOVATE inadvertently reveals the shared name spaces by identifying more appropriate names in natural language to describe the visual contents. This property can be potentially used for analyzing the differences between datasets, merging datasets to construct new benchmarks, and various other data curation processes.
@@
Figure E.4: RENOVATE uncovers segments with shared semantic concepts across datasets.
Appendix F License for Existing Assets